OpenStudy (anonymous):

an errand boy went to the bank to deposit some bills for his employer. Some if the bills were one-dollar bills and the rest were five-dollar bills. The total value of the bills was 86$ and the number of bills was 38. Find the number of each kind of bill. system of equations. Please help if you can thanks!

OpenStudy (anonymous):

Ok so let the number of 1 dollar bills be X and 5 dollar bills be Y. Then sum of bills = 38 i.e. X + Y = 38. And VALUE of bills = 86. i.e. 5Y dollars + X dollars = 86. Hence X+5Y = 86 and X + Y = 38. solve accordingly. :)